Let sin(A) = 3/5 with A in QII. What is the exact value of sin(2A) using an identity.
Answer by lwsshak3(11628)

Let sin(A) = 3/5 with A in QII. What is the exact value of sin(2A) using an identity.
sinA=3/5 (you are working with a 3-4-5 right triangle)
cosA=-4/5
Identity:
sin2A=2sinAcosA=2*(3/5)*(-4/5)=-24/25
..
check: with calculator
A=arcsin(3/5)=143.13º (in QII)
2A=286.26º
sin2A=sin(286.26)≈-0.96
-24/25=-0.96